OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = - 5 x - y = -2 x - y = -1 x - y = - 6

sam (.sam.):

Find the slope

OpenStudy (anonymous):

(x+2)3-7/-2-2=y-3 (x+2)-4/-4=y-3 x+2=y-3 x-y=-5